OpenStudy (anonymous):

a 1 2 - + - + - 6 4 3 I know the lcm is 25, but i do not know where to go from there.

OpenStudy (anonymous):

24*

OpenStudy (agl202):

Are u adding fractions?

OpenStudy (anonymous):

yes, sorry if it was confusing.

OpenStudy (agl202):

Ok, give me sec...

OpenStudy (agl202):

A/6 u have to find a?

OpenStudy (anonymous):

yes, i have to add these together.

OpenStudy (anonymous):

1- find the common denominator

OpenStudy (anonymous):

24

OpenStudy (anonymous):

yes and 12 also

OpenStudy (anonymous):

oh yes, okay so it's 12.

OpenStudy (anonymous):

multiply fractions by a factor to get 12

OpenStudy (anonymous):

so 6x2=12, 4x3=12 3x4=12

OpenStudy (anonymous):

yes but both denominator and nominator

OpenStudy (anonymous):

a/6*2/2 +.....

OpenStudy (anonymous):

so would ax12 just be 12a?

OpenStudy (anonymous):

no a/6*2/2=2a/12

OpenStudy (anonymous):



OpenStudy (anonymous):

so then 1x3=3 and 4x3=12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do we reduce 3/12? so 1/4?

OpenStudy (anonymous):



OpenStudy (anonymous):

then a2/12+3/12+8/12=(2a+11)/12

OpenStudy (anonymous):

OHHH, wait do i include the ( ) in my answer?

OpenStudy (anonymous):

oh nevermind. thank you @AJ01

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

\[\frac{ 2a+11 }{12}\]